OpenStudy (anonymous):

am had a full 1 L bottle of olive oil. He poured 0.25 L of the olive oil into a container to make salad dressing. He also poured an unknown amount x onto a plate as a bread dip. There was 0.6 L of olive oil remaining in the bottle. Which equation can be used to solve for x?

OpenStudy (anonymous):

A. 1 + 0.25 – x = 0.6 B. 1 + 0.25 + x = 0.6 C. 1 – 0.25 – x = 0.6 D. 1 – 0.25 + x = 0.6

OpenStudy (misty1212):

when in doubt, Charlie out

OpenStudy (misty1212):

one minus .25 - x \[1-.25-x=.6\]